OpenStudy (anonymous):

Tonya collected 140 different types of coins from different countries. Fifty of those coins were from Europe. What percent of the coins were from Europe? Round to the nearest percent, if necessary.

OpenStudy (anonymous):

50 / 140 Divide that.

OpenStudy (anonymous):

Then multiply by 100, that's your answer.

OpenStudy (anonymous):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @chel

OpenStudy (mathstudent55):

A percent problem can be solved with a proportion. The total number of coins is 100% of the coins and is 140. The part is 50 coins, and we don't know its percent. let's call the part x percent. The proportion is: 140 is to 100% as 50 is to x%.

OpenStudy (mathstudent55):

Now we write the proportion using fractions: \(\dfrac{140}{100} = \dfrac{50}{x} \) Now you can use cross multiplication to solve for x.

OpenStudy (mathstudent55):

To cross multiply, you multiply diagonally: \(140x = 100 \times 50\) \(140x = 5000\) Now divide both sides by 140 to find x, the percentage of coins from Europe.